consider a line whose slope is 5 and which passes through the point (-2,8) and write the equation of this line in point slope form and slope intercept form




Answer by math_helper(2461)

Point-slope form:
y-8 = 5(x-(-2))
y-8 = 5(x+2)

Slope-intercept form can be found by solving the above for y (and simplifying so it has the form: y = mx + b).